OpenStudy (anonymous):

Tickets to a concert cost $2 for children and $5 for adults. If 630 people attended the concert and the ticket sales totaled $2,910, how many children attended the concert? 36 50 80 120

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

honestly have no idea

OpenStudy (anonymous):

a = amount of adults c = amount of children 5 bucks per adult, sales for that will be 5 * a, or 5a 2 bucks per child, sales for that will be 2 * c, or 2c so 630 people attended, adults and children summed up will be a+c=630 so total sales was $2,910 so adult sales and children sales summed up will be 5a+2c=2910 ------------------------------------------------ from a + c = 630 we could solve for "a" and say that a=630−c so now we know that a = 630-c, so let us use that in the 2nd equation 5(630−c)+2c=2910 solve for "c" to find out how many children attended

OpenStudy (anonymous):

there is help^

OpenStudy (anonymous):

how do i solve for c

OpenStudy (anonymous):

okay if you need help it is 80 that is what the problum says if you solve it it will =80 at the end of you problume

OpenStudy (anonymous):

2 bucks per child, sales for that will be 2 * c, or 2c so 630 people attended, adults and children summed up will be a+c=630 so total sales was $2,910 so adult sales and children sales summed up will be 5a+2c=2910

OpenStudy (anonymous):

okay thank you can you help me on 3 more problems please

OpenStudy (anonymous):

and so ends up 80

OpenStudy (anonymous):

and yes I can

OpenStudy (anonymous):

Tom's and Tim's ages add up to 21 years. In 3 years, Tom will be twice Tim’s age. How old are they now? Tom is 15, and Tim is 6. Tom is 12, and Tim is 9. Tom is 16, and Tim is 5. Tom is 14, and Tim is 7.

OpenStudy (anonymous):

Those tickets are the cheapest concert tickets I've ever heard off.

OpenStudy (anonymous):

open them in a new quistion and medal please

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

@wesleysipes medal please

OpenStudy (anonymous):

@wesleysipes